# Concept-Guided Interpretability via Neural Chunking

### v1

@@ 4.2 Reflection Hypothesis Evaluated on Large Language Models
Figure 3a shows normalized activations of the first 50 neurons (unsorted) across all LLaMA3’s layers for the word “cheese”. Consistent with the reflection hypothesis, the two differently tokenized occurrences of “cheese” evoke similar neural activity patterns, more so than a different word like “ingredient,” suggesting the presence of common neural subpopulation activities encoding recurring concepts. Using recurring words in language as the pattern of interest $s$, we applied PA to identify the neural subpopulation $C(s)$, the mean chunk activity $\overline{\mathbf{h}_{C(s)}}$, and the fluctuation radius $\Delta(s)$ at the last token of each word bearing a variety of tokenization forms.
@@
We generalized our method for identifying concept-encoding neural population chunks to other large-scale models with distinct architectures and sequence-processing mechanisms, including the encoder-decoder model T5 (t5-small ni2021sentence ), the RNN-based RWKV (rwkv-4-169m-pile) peng2023rwkv , and the state-space model Mamba (mamba-130m-hf gu2023mamba ; see Figures 25, 26, 27 in the SI for decoding the example concepts). We evaluated decoding accuracy across these models by testing whether the presence or absence of a concept word could be predicted from chunk activation, going beyond the illustrated concepts to the top 100 most frequent English words. This approach reliably detected concept presence, concept memory, and concept predictions. Additional results are shown in SI C.2. We also confirmed PA’s effectiveness in simpler RNNs (Figure9). Overall, we observed consistent alignment of chunk activations with meaningful input concepts using this method, suggesting that PA yields consistent results across architectures.
Using PA to detect abstract schema and narrative structures
We evaluated PA’s ability to detect more abstract narrative structure by training on 20 stories following a schema (e.g., visit food location → buy item → eat → react) and 13 control stories with narrative structure inconsistent with the schema (examples shown in Figure 5). We extracted shared subpopulation chunks at the end-of-sentence token using PA, and tested on 18 new schema-consistent and 15 inconsistent stories. Schema-related chunks are activated at sentence boundaries in structured narratives but not in controls, suggesting that PA captures abstract narrative schemas beyond surface-level concepts or syntax.
@@
This way, we can simplify high-dimensional embeddings by representing them as the appearance and disappearance of identifiable entities. This enables visualization of the interactions between input tokens and the activations of these recurring entities. Figure 7a illustrates chunk interactions across LLaMA’s layers during early text processing. Recurring chunk patterns emerge, such as similar activations for commas and adjectives in early layers (Figure 7b), which later diverge into distinct neural activations. The full plot is shown in SI D.3. We also interpret the unsupervised learned chunks by correlating them with part-of-speech (POS) tags, the result can be found in SI D.4.
5 Discussion
We identified recurring chunks within the neural dynamics of both RNNs and LLMs, suggesting that the cognitive principle of chunking offers a powerful lens for segmenting high-dimensional neural activity into structured trajectories. This reveals how artificial systems reflect, conceptualize, and abstract patterns in the world. Building on this insight, we propose concept-guided interpretability: a framework that distills complex neural computations into interactions between these recurring entities. Possible future work directions include to further formalize the theoretical foundations, investigate the learning dynamics that give rise to these entities, and leverage this perspective to build a mechanistic understanding of the inner workings of complex learning systems.
Appendix A RNN
A.1 Recurrent Neural Network (RNN) Architecture
The RNN can be described by the following equation:
@@ A.1 Recurrent Neural Network (RNN) Architecture
The hidden state is initialized as all zeros $\mathbf{h}_{0}=\mathbf{0}$
$\mathbf{x}_{n},\mathbf{b}_{h}\in\mathbb{R}^{|\Omega|}$, $\mathbf{h}_{n},\mathbf{b}_{o}\in\mathbb{R}^{d}$, $\mathbf{W}_{ch}\in\mathbb{R}^{d\times(d+|\Omega|)}$, $\mathbf{W}_{co}\in\mathbb{R}^{|\Omega|\times(d+|\Omega|)}$
[…]
@@
To evaluate cross-model consistency, we applied the unsupervised chunk discovery method to RNNs with different architectures. All models exhibited consistent structural regularities (Lf9), with similar chunk counts and alignment to input patterns (Lf12) despite having different architectures.
A.2 Example of a Lookup Table that maps from symbolic clustered state to the input
@@
We apply Discrete Sequence Chunking to the hidden states of a simple RNN which was trained on a sequence of ABCD within default E characters. In table 1, we show the resulting lookup table for the discrete strings corresponding to input signals. Many distinct population states correspond to Es, and for A, B, C, and D, separately, there is a distinct population state for each. Presumably, the network creates additional states also to distinguish Es in different contexts.
A.3 Validation of PA and UCD on simple RNNs
We also validated that PA works for simpler cases on RNNs demonstrated earlier in the paper by applying PA and UCD to RNNs trained on sequences with repeated "ABCD" patterns. PA revealed recurring chunk states whose activation frequency perfectly matched pattern occurrences in the sequence 8. We look at the neural population activity when the network learns from a sequence with repeating ABCD as a cohesive chunk amid background noise composed of random occurrences of E, F, and G. We get the population template by averaging the population activity vector $\mathbf{H}\in\mathbb{R}^{d\times 4}$ over the occurrences of the chunk ABCD:
@@ A.3 Validation of PA and UCD on simple RNNs
Figure 8: Applying population averaging on RNNs predicting simple sequences. Left: The chunk extracted from population averaging align closely to the signal-triggered neural embeddings. Middle: When ABCD does not occur in the sequence, embedding deviates strongly from the extracted chunks. The count of chunk occurrence in the neural population (x axis) correlate with the number of ABCD occurrences in the sequence. Blue line: perfect alignment.
@@
A.4 Pseudocode of Learning Chunks using DSC
Input : $K$, $freq\_threshold$, $symbolized\_neural\_states$, $n\_iter$
Output : $state\_parse$, $vocab$
1 $vocab$leftarrow unique($symbolized\_neural\_states$);
2
[…]
4
[…]
6
7for $n\_iter$ do
8       $ChunkCandidates\leftarrow$ MostCommon($K$,$\text{Counter}(\text{zip}(state\_parse[:-1],state\_parse[1:]))$);
9
10      $merged\_dict\leftarrow$ Merge($c_{L}$, $c_{R}$) $\in$ $ChunkCandidates$ if (count($c_{L}$, $c_{R}$) $\geq$ $freq\_threshold$ $and$ $c_{L}\neq null\_state$ $c_{R}\neq null\_state$);
11
12      $vocab.\text{update}(merged\_dict)$;
13
[…]
15
16      $state\_parse$, $vocab$leftarrow ParseStateSeq($symbolized\_neural\_states$,$vocab$);
17
18
19return $state\_parse$, $vocab$;
Algorithm 1 LearnChunks
1
Input : $trainingsequence$
Output : $state\_parse$, $vocab$
2
3Initialize RNN;
4
[…]
7
[…]
Algorithm 2 Neural Population State Chunking
A.5 The reflection hypothesis on a range of RNN architectures
We show that the reflection hypothesis stays intact independent of RNN’s size. Figure 10 plots RNN’s neural activity with an increasing number of hidden neurons (12, 24, and 48, separately). Across all RNNs, the neural activities reflect regularities in the training sequence.
To evaluate cross-model consistency, we applied the unsupervised chunk discovery method to RNNs with different architectures. UCD learns similar chunk counts and alignment to input patterns (Figure 11) from models with different architectures.
@@ A.5 The reflection hypothesis on a range of RNN architectures
Figure 11: The number of chunks discovered and
A.6 Comparing Trained versus Untrained RNNs
We present a more comprehensive comparison between the trained and untrained RNNs and the characteristics of their neural trajectory patterns as indicated by the chunking method.
@@ A.6 Comparing Trained versus Untrained RNNs
Appendix B Grafting neural chunks versus grafting input
Only chunk-level grafting leads to reliable reuse 13, as Simply grafting the input is ineffective, and just by grafting the learned chunks encourages compositional reuse and improves transfer.
@@ Appendix B Grafting neural chunks versus grafting input
Appendix C Population Averaging
C.1 Layer-wise statistics of example concepts, memory and prediction extracted by PA
Figure 14, 15, 16, and 17 show layer-wise statistics of the learned parameters from the training data, using the population averaging method. From left to right are the optimal tolerance threshold across layers, the number of neurons accountable for the recurring embedding activity, and the maximal deviation threshold.
Figure 14: Example of layer-wise statistics of subpopulation encoding the next step prediction of “cake”.
Figure 15: Layer-wise statistics of subpopulation encoding the next step prediction of the word “in”.
Figure 16: Layer-wise statistics of subpopulation encoding the previous word being “of”.
Figure 17: Example of layer-wise statistics of subpopulation encoding the latest word being “people”.
C.2 Sample Dictionary of Chunks Memorizing and Predicting Signals
We used the population averaging method to extract neural subpopulation activity chunks accounting for the 100 most frequently occurring words in English (Figure 6), adapting the threshold parameters on hidden activities collected from a training prompt and evaluating hidden activities collected from a test prompt (appendix G). For each word, we study subpopulation activities accounting for two steps prior to and two steps subsequent to the last sequence token position. We observed subpopulation chunks that represent information in both directions. Generally, it seems to be the case that the population averaging method learns chunks that are more predictive of the signal in the input sequence with words that have specific meanings than words that serve as prepositions; this can be caused by the network creating many chunks to distinguish a preposition in different contexts, similar to the observation with RNNs (population averaging would fail in this case). We included more examples of chunk evaluation of this type.
Figure 20 visualizes the extracted neural subpopulation for the words “cheese”, “cake”, and “cheesecake”, respectively, in both directions.
We also show more examples of decoding accuracy to evaluate how indicative the extracted sub-population chunk is of the signal’s existence. Note that decoding performance in the 0th layer is usually perfect, as the 0th layer is the embedding layer having a fixed dictionary mapping tokens to the token embeddings. In subsequent layers, the presence of the extract indicates the occurrence of the recurring tokens in the input sequence. Having the recurring token in the input at the last sequence location elicits the network’s hidden activity to lie within the range of the extracted embedding chunk.

### v2

@@ 4.2 Reflection Hypothesis Evaluated on Large Language Models
Figure 2d shows normalized activations of the first 50 neurons (unsorted) across all LLaMA3’s layers for the word “cheese”. Consistent with the reflection hypothesis, the two differently tokenized occurrences of “cheese” evoke similar neural activity patterns, more so than a different word like “ingredient,” suggesting the presence of common neural subpopulation activities encoding recurring concepts. Using recurring words in language as the pattern of interest $s$, we applied PA to identify the neural subpopulation $C(s)$, the mean chunk activity $\overline{\mathbf{h}_{C(s)}}$, and the fluctuation radius $\Delta(s)$ at the last token of each word bearing a variety of tokenization forms.
@@
We generalized our method for identifying concept-encoding neural population chunks to other large-scale models with distinct architectures and sequence-processing mechanisms, including the encoder-decoder model T5 (t5-small ni2021sentence ), the RNN-based RWKV (rwkv-4-169m-pile) peng2023rwkv , and the state-space model Mamba (mamba-130m-hf gu2023mamba ; see Figures 25, 26, 27 in the SI for decoding the example concepts). We evaluated chunk quality by testing whether concepts could be predicted from chunk detections on held-out neural recordings corresponding to test prompts. Beyond the illustrated concepts, we extended the concepts to the top 100 most frequent English words. Overall, we found that PA consistently extracts chunks whose activations correspond to meaningful concepts across diverse architectures. The presence of a chunk reliably predicts the presence of its associated concept in the text with minimal false positives.
Chunks Encoding Abstract Schema and Narrative Structures
We evaluated PA’s ability to find chunks encoding more abstract narrative structure by training on 20 stories following a schema (e.g., visit food location → buy item → eat → react) and 13 control stories with narrative structure inconsistent with the schema (examples shown in Figure 5). We extracted shared subpopulation chunks at the end-of-sentence token using PA, and tested on 18 new schema-consistent and 15 inconsistent stories. Schema-related chunks are activated at sentence boundaries in structured narratives but not in controls, suggesting that PA captures abstract narrative schemas beyond surface-level concepts or syntax.
@@
We then evaluate the general effectiveness of grafting general concepts on the ROCStories benchmark mostafazadeh2016corpus , sampling 2,000 sentences to extract neural population chunks corresponding to the top 20 frequent concept words (filtered to avoid short or overlapping terms) in addition to the existing concepts collected. To assess context-dependent effects of prompt used when grafting the concept-encoding neural chunks, we used prompts from the TREC Question Classification dataset li2002learning , which includes six coarse-grained context categories. For each category, we sampled 50 prompts and evaluated grafting success by measuring the difference in concept occurrence probability between grafted and ungrafted (control) LLaMA-3 generations. To reduce computational cost, we categorized grafting into early (layers 1–9), middle (10–19), and late (20–29) interventions. Results are shown in Table 1. We found that grafting consistently increased concept occurrence across contexts, with early-layer grafting being the most effective.
Table 1: Effect of grafting on TREC categories (percentages).
TREC Category No Graf. Early Graf. Middle Graf. Late Graf.
ABBR (Abbreviations, acronyms) 14.9% 55.9% 30.8% 18.0%
DESC (Descriptions, definitions) 15.6% 49.0% 28.1% 20.7%
ENTY (Entities) 12.6% 48.1% 22.5% 16.9%
HUM (Human-related) 11.9% 46.7% 21.5% 15.2%
LOC (Locations) 10.7% 47.5% 20.5% 14.4%
NUM (Numeric answers) 11.5% 45.3% 21.8% 16.0%
@@
This way, we can simplify high-dimensional embeddings by representing them as the appearance and disappearance of identifiable chunks. This enables visualization of the interactions between input tokens and the activations of these recurring entities. Figure 6a illustrates chunk interactions across LLaMA’s layers during early text processing. Recurring chunk patterns emerge. For example, commas and adjectives show similar activations in early layers (Figure 6b), but their representations later elicit distinct chunks in deeper layers. The full plot can be found shown in SI F.5. We also interpret the unsupervised learned chunks by correlating them with part-of-speech (POS) tags, the result can be found in SI F.6.
5 Conclusion, Discussion
We propose the Reflection Hypothesis, which posits that neural population activity reflects the structure of the data. We provide evidence in support of this hypothesis and, building on this insight, leverage the cognitive principle of chunking to isolate perceptual entities as units of interpretation. Using the methods we introduce, we identify chunks and test their feasibility across models ranging from small-scale RNNs to a number of large-scale LLMs. We find that these chunks correlate strongly with both concrete and abstract concepts—spanning recurring words, sentence structure, and POS tags. These results suggest that chunking provides a powerful lens for segmenting high-dimensional neural activity into structured trajectories. This work invites future research to formalize its theoretical foundations, investigate the learning dynamics that give rise to such entities, and apply this perspective to build a mechanistic understanding of the inner workings of complex learning systems.
Supplementary Information (SI)
Appendix A Extended Discussion on Related Work
A.1 Comparison with Other Interpretability Approaches
Our work is a different approach from a number of other interpretability methods. For example, Black et al. define polytopes as linear regions in input space and study their geometry [12]. Vielhaben et al. define concepts as linear subspaces in feature space, spanned by an unbounded number of basis vectors [83]. Our definition of a chunk is compact and bounded subspace, hence it does not tolerate addition or scalar multiplication of some basis vectors. In fact we would expect a certain level of scalar multiplication may run into other chunks. We suggest that chunks themselves are interpretable.
Other approaches, such as Tamkin et al., propose architectures that enforce sparsity and interpretability. In particular, they learn a codebook of features through codebook bottlenecks [78], which requires training new models with modified architectures. By contrast, our goal is to interpret the raw activity of large models directly, independent of their training procedures or architectures, rather than optimizing a new model.
A.2 Comparison with SAEs and Representation Engineering
Our method for grafting and freezing corresponds to steering and unlearning in SAE work [29, 57]. However, the chunk grafting is simpler and more direct than that of SAEs. In SAEs, steering is typically achieved by clamping a feature to a specific value—either through a constant multiplier or a scaled value relative to its maximum activation [81]. This clamped feature then probabilistically influences the network’s output during text generation.
However, SAE features often do not correspond clearly to human-interpretable concepts. The mapping between features and words is indirect, requiring additional steps such as VocabProj or MaxAct to identify associated tokens [29]. This makes steering both complex and fragile. Interpreting SAE features remains an open challenge, and developing effective steering methods is still an active area of research [77].
In contrast, PA offers more interpretable units by design. Chunks are derived from recurring patterns in the model’s activations, grounded in actual input semantics. This enables more transparent and effective interventions, such as concept grafting or unlearning, without the need for specialized heuristics. We also evaluated our steering efficiency on a bigger corpus across multiple contexts, and verified that the extracted chunks can be an efficient way of steering network behavior.
The other main difference between this approach and SAE is compute efficiency. The population averaging (PA) method is a post hoc analysis that requires no optimization or gradient computation. Neural activations are precomputed, and the only cost lies in computing prototype vectors and evaluating distances across a small number of deviation thresholds—operations that scale linearly with the number of tokens and the embedding dimension.
In contrast, sparse autoencoders (SAEs) require multiple forward and backward passes through both encoder and decoder networks, typically over many training epochs. This results in orders-of-magnitude greater computational cost. PA, by contrast, offers a lightweight and accessible alternative for large-scale interpretability research. The unsupervised chunk discovery method also remains computationally feasible. It runs efficiently on a single GPU. Given precomputed neural activations, it can be applied broadly to various architectures and datasets.
This work also differs from representation engineering, which isolates neural activity associated with a predefined concept or function, such as bias or truthfulness [94]. Representation engineering typically fits a linear model to identify activity directions that predict or intervene on the target concept. However, this approach assumes the concept of interest is known a priori. In contrast, we assign a single chunk to summarize the neural activity of a concept, enabling discovery without prior labels. In contrast to grafting with chunks, which directly overwrites neural activity with a given chunk, representation engineering applies a latent direction whose effect depends on a chosen magnitude. Determining this strength is a nontrivial hyperparameter choice and not always straightforward.
A.3 Comparison with Concepted-based Interpretability Literature on Vision Models
Our work is different from concept-based interpretability methods for image classifiers. In [34], [30] and [45] concepts correspond to visual features in the image dataset (e.g. pixels of a wheel), and they study what part of the image influences the network’s classification decision. In contrast, our work primarily studies network interpretability processing sequences of tokens. Our work examines how the internal structure of the network exhibits recurring patterns of entities. We reframe interpretability by connecting neural population trajectories to human chunking.
Appendix B Definition of Grafting and Freezing
Grafting refers to the targeted insertion of previously learned chunks into a neural network’s internal computation. We modify its internal activations $\mathbf{h}$ at a specified layer $l$ and neural subset $C$ by grafting the chunk $\mathbf{h}_{C,l}\leftarrow\overline{\mathbf{h}_{C,l}}$ to replace the original neural subpopulation activities.
Freezing refers to the operation of zeroing out the neural activations that support the chunk. This is implemented by replacing the network activation with stored chunk representations at the corresponding neurons. At a given layer $l$, we enforce: $\mathbf{h}_{C,l}=\mathbf{0}$.
Appendix C Pseudocode of Learning Chunks Using Discrete Sequence Chunking
Input : $K$, freq_thresholdfreq\_threshold, symbolized_neural_statessymbolized\_neural\_states, n_itern\_iter
Output : state_parsestate\_parse, vocabvocab
1 vocabvocab $\leftarrow$ unique(symbolized_neural_statessymbolized\_neural\_states);
2
3null_statenull\_state $\leftarrow$ GetMostFreq(symbolized_neural_statessymbolized\_neural\_states);
4
[…]
6
7for n_itern\_iter do
[…]
9
10 $merged\_dict\leftarrow$ Merge($c_{L}$, $c_{R}$) $\in$ ChunkCandidatesChunkCandidates if (count($c_{L}$, $c_{R}$) $\geq$ freq_thresholdfreq\_threshold andand $c_{L}\neq null\_state$ $c_{R}\neq null\_state$);
11
12 $vocab.\text{update}(merged\_dict)$;
13
14 $\text{MergeOverlappingChunks}(vocab)$;
15
16 state_parsestate\_parse, vocabvocab $\leftarrow$ ParseStateSeq(symbolized_neural_statessymbolized\_neural\_states,vocabvocab);
17
18
19return state_parsestate\_parse, vocabvocab;
Algorithm 1 LearnChunks
1
Input : trainingsequencetrainingsequence
Output : state_parsestate\_parse, vocabvocab
2
3Initialize RNN;
4
5hidden_stateshidden\_states, sequencesequence $\leftarrow$ TrainRNN(sequencesequence)
6symbolic_hidden_statessymbolic\_hidden\_states $\leftarrow$ ClusterAndAssignSymbolToEachNeuron(hidden_stateshidden\_states);
7
8state_parsestate\_parse, vocabvocab $\leftarrow$ LearnChunks(symbolic_hidden_activitysymbolic\_hidden\_activity, sequencesequence);
Algorithm 2 Neural Population State Chunking
Appendix D RNN
D.1 Recurrent Neural Network (RNN) Architecture
We provide more details on the structure of RNN. The RNN used by this work can be described by the following equation:
@@ D.1 Recurrent Neural Network (RNN) Architecture
The hidden state is initialized as all zeros $\mathbf{h}_{0}=\mathbf{0}$. $\mathbf{x}_{n},\mathbf{b}_{h}\in\mathbb{R}^{|\Omega|}$, $\mathbf{h}_{n},\mathbf{b}_{o}\in\mathbb{R}^{d}$, $\mathbf{W}_{ch}\in\mathbb{R}^{d\times(d+|\Omega|)}$, $\mathbf{W}_{co}\in\mathbb{R}^{|\Omega|\times(d+|\Omega|)}$.
We implemented a simple RNN with 12 hidden units, consisting of two linear layers: one for updating the hidden state and another for generating output predictions. At each time step, the input vector is concatenated with the previous hidden state and passed through these layers. The output is normalized using log softmax to produce a probability distribution over the output classes. The RNN is optimized with cross-entropy loss using Adam ($\text{learning rate}=0.005$). Training is conducted on random subsequences of length 200 per batch, with the hidden state initialized to zero at the start of training. We track hidden states for analysis.
@@
To evaluate cross-model consistency, we applied the unsupervised chunk discovery method to RNNs with different architectures. All models exhibited consistent structural regularities (Figure 10), with similar chunk counts and alignment to input patterns (Figure 11) despite having different architectures.
D.2 Example of a Lookup Table that maps from symbolic clustered state to the input
@@
We apply Discrete Sequence Chunking to the hidden states of a simple RNN which was trained on a sequence of ABCD within default E characters. In table˜2, we show the resulting lookup table for the discrete strings corresponding to input signals. Many distinct population states correspond to Es, and for A, B, C, and D, separately, there is a distinct population state for each. Presumably, the network creates additional states also to distinguish Es in different contexts.
D.3 Validation of PA on RNNs
We also validated that PA works for simpler cases on RNNs demonstrated earlier in the paper by applying PA and UCD to RNNs trained on sequences with repeated "ABCD" patterns. PA revealed recurring chunk states whose activation frequency perfectly matched pattern occurrences in the sequence 8.
We look at the neural population activity when the network learns from a sequence with repeating ABCD as a cohesive chunk amid background noise composed of random occurrences of E, F, and G. We get the population template by averaging the population activity vector $\mathbf{H}\in\mathbb{R}^{d\times 4}$ over the occurrences of the chunk ABCD:
@@ D.3 Validation of PA on RNNs
Figure 8: Applying population averaging on RNNs predicting simple sequences. Left: The chunk extracted from population averaging aligns closely with the signal-triggered neural embeddings. Middle: When ABCD does not occur in the sequence, embedding deviates strongly from the extracted chunks. The count of chunk occurrence in the neural population (x axis) correlates with the number of ABCD occurrences in the sequence. Blue line: perfect alignment.
@@
D.4 Validate UCD on RNNs
To evaluate cross-model consistency, we applied the unsupervised chunk discovery method to the simple example in RNN. UCD learns similar chunks which aligns with the DSC extracted chunks (Figure 9).
D.5 The Reflection Hypothesis Across RNN Architectures
We show that the reflection hypothesis stays intact, independent of the RNN’s size. Figure 10 plots RNN’s neural activity with an increasing number of hidden neurons (12, 24, and 48, separately). Across all RNNs, the neural activities reflect regularities in the training sequence.
Moreover, applying DSC on the RNNs suggests that the number of chunks extracted stays consistent across RNNs with different numbers of hidden units (Figure 11).
@@ D.5 The Reflection Hypothesis Across RNN Architectures
Figure 11: The number of chunks stays consistent across RNN with different number of hidden units.
D.6 Contrasting Trained versus Untrained RNNs
We provide a comprehensive comparison of trained and untrained RNNs, focusing on the characteristics of their neural trajectory patterns as revealed by the chunking method.
@@ D.6 Contrasting Trained versus Untrained RNNs
D.7 Grafting Neural Chunks versus Grafting Input
We examine the effects of grafting at two levels—neural population chunks and input—on the behavior of RNNs when transferring to sequences with compositional components. Chunk-level grafting leads to reliable reuse (Fig. 13); input grafting proves ineffective. This suggests that merely changing the input is insufficient to alter network behavior; rather, behavior is shaped by the coordinated activity of a larger set of computing units beyond those directly encoding the input.
@@ D.7 Grafting Neural Chunks versus Grafting Input
Appendix E Population Averaging
E.1 Example Layer-wise Statistics PA-extracted Chunks
Beyond using PA to identify concept-encoding chunks for concepts occurring at the end of a prompt sequence, the same approach can also be applied to other positions in the sequence. Specifically, PA can isolate chunks that encode concepts appearing at earlier points in the sequence (prior to the end token), as well as chunks that encode the future occurrence of a target concept. The latter correspond to invariant neural subpopulations whose activity reliably precedes the later appearance of a given concept, which corresponds to predictive structure in the network’s dynamics.
Figures 14, 15, 16, and 17 present layer-wise statistics of the learned parameters obtained from the training data using the population averaging (PA) method. From left to right, the plots display: Left, the optimal tolerance threshold across layers, Middle the number of neurons contributing to recurring embedding activity, and Right the maximal deviation threshold.
Note that the maximal deviation learned by the method typically increases from earlier to later layers. This trend reflects the architecture of LLaMA, where deeper layers exhibit a broader range of variability in their neural activity. In addition, the analysis reveals that concept encoding is generally supported by substantial populations of neurons rather than isolated units. Sparse encodings of concepts are rare. This suggest that concepts emerge through distributed patterns of activity across the network rather than through localized neural encoding.
Figure 14: Layer-wise statistics of PA extracted chunks commonly preceding the concept “cake”.
Figure 15: Layer-wise statistics of PA extracted chunks commonly preceding the concept “in”.
Figure 16: Layer-wise statistics of PA extracted chunks commonly proceeding the concept “of”.
Figure 17: Layer-wise statistics of PA extracted chunks encoding the latest word being “people”.
E.2 Sample Dictionary of Chunks Memorizing and Predicting Signals in Text